OpenStudy (anonymous):

How is: sqrt(a^2-a^2sin^2(x) =a*sqrt(1-sin^2(x))

OpenStudy (anonymous):

\[\sqrt{(a^2-a^2\sin^2(x)} =a\sqrt{(1-\sin^2(x))}\]

OpenStudy (anonymous):

sqrt(a^2) = a.

OpenStudy (anonymous):